A particle travels in a circle of radius 20 cm at a speed that uniformly increases. If the speed changes from 5 m/s to 6 m/s in 2 s, find angular acceleration.

Grade:12th pass

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
The tangential acceleration is given by

a_{t} = \frac{\mathrm{d} v}{\mathrm{d} t}
= \frac{6-5}{2} m/s^{2} = 0.5 m/s^{2} .
The angular acceleration isa = \frac{a_{t}}{r}
\frac{0.5 m/s^{2}}{20 cm} = 2.5 rad/s^{2}